OpenStudy (anonymous):

Simplify 6√44-√45-√99+√20

OpenStudy (anonymous):

-6√80 is what I got

OpenStudy (anonymous):

@MathHelper22 how did u get this answer? can u pls show us your work?

OpenStudy (anonymous):

if its not right im not gonan waste my time?

OpenStudy (anonymous):

seems ur answer is wrong but would like to see your method first so that ur mistake can be rectified

OpenStudy (mertsj):

\[6\sqrt{44}-\sqrt{45}-\sqrt{99}+\sqrt{20}=6\sqrt{4(11)}-\sqrt{9(5)}-\sqrt{9(11)}+\sqrt{4(5)}\]

OpenStudy (mertsj):

\[6\sqrt{4}\sqrt{11}-\sqrt{9}\sqrt{5}-\sqrt{9}\sqrt{11}+\sqrt{4}\sqrt{5}\]

OpenStudy (mertsj):

Perhaps that will help.

OpenStudy (anonymous):

got it its 9√11-√5

OpenStudy (mertsj):

yep